Solve each system of inequalities by graphing
x+8/3y=12
1/2x+4/3y=5
I tried x+8/3y=12 .5x+4/3y=6
-x -x -.5x -.5x
3(8/3)y=(-x+12)3 3(4/3y)=(-.5x+6)3
8y=-3/8x+36/8 4y=-1.5x+18
y=-3/8x+4.5 /4 /4
y=-1.5/4x+4.5

Answer by stanbon(75887)

Solve each system of inequalities by graphing.
Comment: These are not inequalities.
x+8/3y=12
1/2x+4/3y=5
--------------
Rearrange each equation to get slope-intercept form.
x+8 = 36y
y = (1/36)x + (2/9)
Plot the points: (0, 2/9) and (36,11/9)
---
3x + 16y = 30
y = (-3/16)x + (15/8)
Plot the points: (0,15/8) and (16,-9/8)
-----------------------------

===============================================
Ans: x = 7.6774.. y = 0.4355.. when graphed with a calculator.
===============================================
Cheers,
Stan H.